Solve x2+4x-7=0 using quadratic equation and then approximate answer to three decimal places using a calculator
Answer by Jolliano(16)

USING COMPLETING THE SQUARE METHOD.
X^2 + 4x - 7 = 0
X^2 + 4x = 7
Find half the coefficient of x,square it and then add it to both sides.
X^2 + 4x + (2^2) = 7 + 2^2
x^2 + 4x + 4 = 7 + 4 Factorise the L.H.S,
(x + 2)^2 = 11
Find the root of both sides,
(x+2) = +√ 11 or -√ 11 Taking positive root first,
x + 2 = 3.317
x = 3.317 - 2
x = 1.317
Taking negative root, x + 2 = -3.317
x = -3.317 - 2
x = -5.317